REACTION PATH HAMILTONIAN CALCULATION OF TUNNELING SPLITTING IN PROTONATED METHANOL AND METHYLAMINE.

\underline{A. R SHARMA}, J. M BOWMAN, Cherry L. Emerson Center for Scientific Computation,Department of Chemistry, Emory University, Atlanta GA 30322, USA; S. C CARTER, Department of Chemistry, University of Reading, RG6 2AD, England.

Both protonated methanol(CH_{3}OH_{2}^{+}) and methylamine(CH_{3}NH_{2}) show two interesting large amplitude motion, namely, internal rotation and inversion. The internal rotation spectrum results from the rotation of the CH_{3} moiety against the OH_{2} group in case of protonated methanol and of NH_{2} group in the case of methylamine. The other interesting phenomenon is inversion involving the OH_{2} group and NH_{2} group in the case of protonated methanol and methylamine, respectively. The influence of inversion in methylamine, on bands in the near and middle infrared has been investigated. Microwave and far-IR spectrum of methylamine has been measured and spectral frequency calculations derived from rotation-internal rotation-inversion analysis has been reported. However, in the past, detection of protonated methanol (CH_{3}OH_{2}^{+}), in interstellar clouds has been impossible due to the lack of both laboratory spectra and calculation.

In this contribution ab initio based potential energy surface (PES) and dipole moment surface (DMS) is presented for both protonated methanol and methylamine. The PES and DMS are developed by least squares fitting of ab initio energy values computed at CCSD(T)/AVTZ level of theory and dipole moments at MP2/AVTZ level of theory. Internal rotation and inversion transition states and normal-mode frequencies will be reported. One-dimensional tunneling splitting calculations will be reported. Tunneling splittings at higher dimensionality, ro-vibrational states and transition intensities are calculated using reaction path Hamiltonian(RPH) as implemented in MULTIMODE(MM) code. MM treats polyatomic molecules with large-amplitude motion as and one special coordinate which is the large-amplitude vibrational coordinate. Complete integration is performed over reaction path coordinate, and the N-mode MULTIMODE coupling approximation for the evaluation of the matrix elements applies only to the 3N - 7 normal coordinates.